DIRC Operations Weekly Report

Period: Wednesday 2007/4/24 - Wednesday 2007/5/02

Oncall: Jose Benitez
Fast Monitoring: Jose Benitez, Alejandro Perez

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070502.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070502

Contents


BaBar-PEP news

PEP peak lumi: ~7.0 x 1033 (previous week: ~6.0 x 1033)

BaBar integrated lumi: 1.8 fb-1 (previous week:2.13 fb-1)

Run 6 PEP delivery: 1.8 fb-1 (previous week: 2.19 fb-1)

Run 6 total delivery: ~32.4 - ~30.61 fb-1 Run 6 total BaBar recorded: 31.2 - 29.38 fb-1 including 0 fb-1 off-peak data

PEP activities this week:

  • 4/26:ROD due to vacuum problem, chamber sealed last time still leaking. Leak could not be found. Chamber to be replaced.
  • 5/1 : Chamber replacement started.
    Problem: Q2 failed to ramp down, Q2 melted and water hose broke when the cooling wast turned off.
    Consequences: All systems powered off. Much water spilled in DIRC tunnel, flowed mostly backward but also forward. Q2 will be replaced => ROD will take ~1week longer.

    BaBar activities/issues this week:

  • 5/1: Water spill got to EMC electronics.
  • 5/3: DCH and DIRC turned back on succesfully.

    DIRC activities